OpenStudy (anonymous):

integral of (2x-3)/(sqrt(4x-(x^2)))

OpenStudy (anonymous):

Complete the square: \[\begin{align*} 4x-x^2&=-(x^2-4x)\\ &=-(x^2-4x+4-4)\\ &=-\left((x-2)^2-4\right)\\ &=4-(x-2)^2 \end{align*}\] \[\int\frac{2x-3}{\sqrt{4x-x^2}}~dx=\int\frac{2x-3}{\sqrt{4-(x-2)^2}}~dx\] Let \(x-2=2\sin u\), then \(dx=2\cos u ~du\). This also gives you \(2x-3=4\sin u+1\). \[\int\frac{(4\sin u+1)(2\cos u)}{\sqrt{4-(2\sin u)^2}}~du\\ \int\frac{(4\sin u+1)\cos u}{\sqrt{1-\sin^2u}}~du\\ \int\frac{(4\sin u+1)\cos u}{\sqrt{\cos^2u}}~du\\ \int(4\sin u+1)~du\]